Please explain that tan theta + cot theta =2 how do we get value of theta?

Grade:10

1 Answers

Defender Singh gurjar
6 years ago
We can write this eq. as tanx+cotx-2=0 Then,Tanx+1/tanx-2=0..........(1)(Tanx)^2+1-2tanx=0This is a identity of (Tanx-1)^2=0From this we getTanx-1=0So,Tanx=1..............(2)NowPutting value of tanx in eq. 1 we getCotx=1.........(3)From the eq. (2) and (3) we get1+1-2=0SoL.H.S= R.H.S.Hence proved